OpenStudy (anonymous):

Is there a difference between a line with a slope of 0 and a line with a undefined slope?

OpenStudy (anonymous):

Recalling the equation for finding slope of a line with two points being m = (y2 - y1) / (x2 - x1) A slope of zero means that y2-y1 has to be zero, meaning that y doesnt change over a change in x, so the line is a straight horizontal line. m = (0)/ (x2 -x1) = 0 An undefined slope means that x2-x1 is equal to zero meaning that x doesnt change over a change in y, meaning that the line is a straight vertical line. m = (y2-y1) / 0 = undefined